# FirstReferences

### v2

@@ Abstract
Figure 4: Transfer rates in the VAET regime. (A) Transfer rate spectrum of degenerate VAET ($\omega_{1}\!=\!\omega_{2}\!\equiv\!\omega$) with $(V,g_{1},\gamma_{1},g_{2},\gamma_{2})\!=\!(0.200,0.200,0.045,0.220,0.018)\omega$. Red circles represent experimental data with error bars estimated via bootstrapping (see Methods). The solid red curve shows the transfer rate calculated from Eq. (2), using the definition in Eq. (3) and including spin decoherence ($\gamma_{z}=0.0007\omega$) and motional dephasing of both modes ($\gamma_{im}=0.0020\omega$, with $i=1,2$). The blue curve shows the numerical result of the single-mode VAET case, where $\omega_{2}\!=\!g_{2}\!=\!\gamma_{2}\!=\!\gamma_{2m}\!=\!0$. (B) Experimental and numerical donor population evolution $P_{D}(t)$ versus the number of vibrational oscillations $\omega t/2\pi$, using the same parameters as in the red circles and solid red curve in (A), respectively. Here, the detunings from the two tilt modes for encoding the degenerate vibrational energies are both set to $\delta_{1}=\delta_{2}=-2\pi\times 10$ kHz. (C and D) The same layout as in (A and B) for the non-degenerate VAET case ($\omega_{1}\!>\!\omega_{2}$) with $(V,g_{1},\gamma_{1},\omega_{2},\gamma_{z},\gamma_{im})=(0.167,0.275,0.028,0.667,0.0006,0.0017)\omega_{1}$ and $(g_{2},\gamma_{2})=(0.263,0.019)\omega_{2}$. The characteristic timescale is defined by the fast vibrational frequency $\omega_{1}$. In this case, the detuning from the inner (outer) tilt mode, which encodes the fast (slow) vibrational energy, is set to $\delta_{1}=-2\pi\times 12$ kHz ($\delta_{2}=-2\pi\times 8$ kHz). The red triangle, orange pentagon, green square, and blue circle markers correspond to $\Delta E=\{1.00,1.20,1.50,1.95\}\omega$ in the degenerate case and $\Delta E=\{0.75,0.92,1.42,1.60\}\omega_{1}$ in the non-degenerate case, respectively. Near each resonant peak in the rate spectra, we label the total vibrational energy involved in the transfer.
@@
Vibrationally assisted exciton transfer (VAET) - Unlike in the CT regime, where strong vibronic coupling distorts the donor-acceptor potential energy landscape that defines the vibronic eigenstates of the system, in the VAET regime ($g_{i}\!\ll\!\omega_{i}$), the vibrational modes are weakly coupled to the electronic degree of freedom and therefore only act as facilitators of the exothermic transfer, where units of vibrational energy are exchanged with the electronic sites for the excitation transfer to occur at $\Delta E\approx\sqrt{(\ell_{1}\omega_{1}+\ell_{2}\omega_{2})^{2}-(2V)^{2}}$ (see Fig. 4) [30]. Here, we also consider systems with strong electronic coupling ($|V|\sim\lambda_{i}/4$), which again places the system in the non-perturbative regime. In the vibrationally degenerate case ($\omega_{1}=\omega_{2}\equiv\omega$) with Fig. 4A, the transfer rate of the first resonance in the two-mode model is not enhanced compared to the single-mode case because only a single phonon from either degenerate mode can contribute to the transfer process at a time. It actually becomes slower because of the additional broadening from the dissipation of the second mode. On the contrary, at $\Delta E\approx\sqrt{(2\omega)^{2}-(2V)^{2}}$, processes in which the total vibrational energy of $2\omega$ is given by a linear combination of the energy quanta of the two degenerate modes ($2\omega_{1}=2\omega_{2}=\omega_{1}+\omega_{2}$) interfere constructively, resulting in enhanced transfer rates compared to the single-mode case, where only the process with the vibrational energy input of $2\omega_{1}$ can assist the transfer.
@@
In the non-degenerate case $(\omega_{1}>\omega_{2})$, these second-order processes occur at different donor-acceptor energy offsets, and their corresponding transfer resonances can be resolved experimentally (see Fig. 4C). As such, the presence of the slow mode does not affect the transfer rates at the resonances involving $\omega_{1}$ and $2\omega_{1}$ energy inputs. This is because, when we choose $\omega_{1}/\omega_{2}$ to be non-integer, there exists no non-trivial resonance involving an energy input of $\ell_{2}\omega_{2}$ from the slow mode, where $\ell_{1}\omega_{1}+\ell_{2}\omega_{2}=\ell\omega_{1}$. However, additional resonances emerge instead from processes involving energy inputs given by linear combinations of the two non-degenerate vibrational energies. In Fig. 4C, we observe three additional resonances beyond the two existing resonances provided by the fast vibrational mode for the two-mode case. These peaks correspond to the processes involving vibrational energy inputs of $\omega_{2}$, $2\omega_{2}$, and $\omega_{1}+\omega_{2}$.
@@
It is worth noting that, in both cases of mode degeneracy, the second-order processes associated with a vibrational energy contribution of $\omega_{1}+\omega_{2}$ consist of two pathways (one phonon from the fast mode, then another from the slow mode, and vice versa) that interfere constructively. This coherent addition of the two rate amplitudes leads to an approximate two-fold enhancement of the peak at $\Delta E\approx 1.63\omega_{1}$ in Fig. 4C with respect to the neighboring peaks at $\Delta E\approx 1.29\omega_{1}$ and $\Delta E\approx 1.97\omega_{1}$ that are associated with the processes involving $2\omega_{2}$ and $2\omega_{1}$ energy inputs, respectively (see Appendix E).
@@
In this work, we leverage the remarkable tunability of the trapped-ion platform to experimentally realize an open two-mode LVCM in two phenomenologically distinct regimes associated with CT and VAET, as well as their crossover. We simultaneously apply twelve carefully calibrated laser tones to independently control the coherent evolution of the qubit and the damping rates of two bosonic modes in the multi-species ion system. We observe enhanced transfer rates due to the presence of the second mode across all vibronic coupling regimes—regardless of mode degeneracy—pointing out their differences and similarities. Our conclusions can be extended to two-site LVCM systems with more than two vibrational modes, which exhibit the same qualitative features identified in our findings (see Appendix F). This observation also highlights the necessity of considering anharmonicity in quantum dynamics models, where quantum scrambling can occur at resonances, transitioning coherent quantum behavior into chaotic dynamics [35].
@@
The experimental toolbox we deploy here is intrinsically scalable, as the same ion-laser couplings and sympathetic cooling techniques used to realize a two–mode, two–site LVCM can be extended to many vibrational modes and electronic sites without introducing additional physical overhead. In trapped-ion hardware, it is possible to include more molecular sites, as each extra qubit ion supplies a fully controllable two-level system that can encode a chromophore or charge-transfer center. Each additional qubit ion or coolant ion also adds three collective bosonic modes, which can be directly used for reservoir engineering by tailoring the spectrum of sympathetic-cooling lasers, without the need for digitization [36]. An arbitrary subset of these modes can then be endowed with individually programmable frequencies, coupling strengths, and dissipation rates, allowing the simulation of dissipative chemical dynamics in complex solvent environments [37, 38]. Employing multiple engineered bosonic modes also enables the experimental realization of spin-boson models with tunable spectral densities, formed by a linear superposition of Lorentzian components [39, 33, 21], crucial for exploring phenomena related to non-Markovian dynamics [40], such as coherence trapping [41, 42, 43, 44], and dissipative quantum state engineering [45].
@@ Methods
The experimental system used for this study is thoroughly described in Ref. [20], where the dynamics of the single-mode LVCM in the CT regime is realized with four Raman 355 nm and two 435.5 nm laser tones. In this work, we include two additional tones to each beam to generate the terms associated with the second vibrational mode, resulting in a total of ten laser tones on the 355 nm and 435.5 nm lasers, over which we have full control of both amplitudes and frequencies. We also use the collective modes of the two-ion chain along both radial directions of the trap—specifically the $y$ and $z$ tilt modes ($\omega_{\text{tilt},y}\equiv\omega_{\text{tilt},1}=2\pi\times 3.151$ MHz and $\omega_{\text{tilt},z}\equiv\omega_{\text{tilt},2}=2\pi\times 3.740$ MHz)—to encode the vibrational degrees of freedom in the two-mode LVCM Hamiltonian. The unused radial collective modes are the center-of-mass modes with $\omega_{\text{com},y}\equiv\omega_{\text{com},1}=2\pi\times 3.318$ MHz and $\omega_{\text{com},z}\equiv\omega_{\text{com},2}=2\pi\times 3.882$ MHz. Since the frequency separations among the available radial collective modes are sufficiently large ($\Delta\omega_{\text{trap},j}\gtrsim 2\pi\times 140$ kHz), undesired off-resonant spin-phonon interactions, which are proportional to $g_{j}/(\mu_{i}-\omega_{\text{trap},j})$, can be neglected [48]. Here, $\mu_{i}$ is the 355 nm laser frequency detuning from the qubit resonance, used to generate the spin-dependent force on the target radial tilt mode $i$, and $\omega_{\text{trap},j}$ is the relevant collective radial frequency considered for the off-resonant spin-phonon drive ($j\neq i$).
@@
The experimental sequence (see Fig. 1C) begins with Doppler cooling and Raman-resolved sideband cooling on all four collective radial modes of the chain, which results in an initial phonon population of both radial tilt modes, $\bar{n}_{0,i}\sim 0.1$–0.2, which is set to match the independently measured $\bar{n}_{i}$. We then apply a $\pi/2$ pulse to map the $z$ qubit basis to the $y$ basis and two consecutive displacement operations via the spin-dependent optical force to prepare the system in the donor vibronic state, |$\ket{D}\bra{D}\otimes\rho_{1-}\otimes\rho_{2-}$, where $\rho_{i-}=\sum_{n_{i}}e^{-n_{i}\omega_{i}/k_{B}T_{i}}\ket{n_{i-}}\bra{n_{i-}}$ | represents a thermal state with temperature $k_{B}T_{i}\approx\omega_{i}/\log(1+1/\bar{n}_{i})$, and |$\ket{n_{i\pm}}=\mathcal{D}(\pm g_{i}/2\omega_{i})\ket{n_{i}}$ are displaced Fock states for vibrational mode $i$. For simulating the open-system LVCM dynamics described by Eq. (2), we simultaneously apply the six 355 nm, four 435.5 nm, and two 935 nm laser tones. After evolving the system for time $t_{\rm sim}$, we apply another $\pi/2$ pulse to map the quantum state in the $y$ basis back to the $z$ qubit basis and measure the probability of the system being in the donor state $P_{D}=(\langle\sigma_{z}\rangle+1)/2$ via state-dependent fluorescence.
@@
We experimentally simulate the LVCM dynamics over a finite duration of $t_{\rm sim}=$ 2–8 ms, corresponding to 24–65 vibrational cycles of the fast mode (defined by $\omega_{1}t/2\pi$), within which the system undergoing the first transfer resonance reaches equilibrium. Therefore, we need to remove the contribution from the finite-time evaluation of the non-zero steady-state donor population in the transfer rate calculations, as follows [20, 34]:
[…]
By interpolating the donor population probability $P_{D}(t)$ for both the experimental and numerical data using the same time steps and applying the modified formula above, we obtain the transfer rates of the dynamics reported in the main text. As in Ref. [20], we also use a resampling (bootstrapping) procedure at each time step of the $P_{D}(t)$ measurements to estimate the uncertainties of the extracted transfer rates. For each time step, we treat the experimental error as the standard deviation of a normal distribution centered at the measured mean value, from which the resampled datasets are drawn. We then use the standard deviation of the transfer rates calculated from the resampled datasets as the uncertainty associated with the reported rates.

### v3

@@ Abstract
Figure 4: Transitioning from CT to VAET ($\omega_{1}\!>\!\omega_{2}$). (A) Transfer rate spectrum for $(V,g_{1},\gamma_{1},\omega_{2})=(0.063,1.288,0.015,0.375)\omega_{1}$ and $(g_{2},\gamma_{2})=(0.607,0.067)\omega_{2}$. Red circles represent experimental data with error bars estimated via bootstrapping (see Methods). The solid red curve shows the transfer rate calculated from Eq. (4), using the definition in Eq. (5) and including spin decoherence ($\gamma_{z}=0.0009\omega_{1}$) and motional dephasing of both modes ($\gamma_{im}=0.0100\omega_{1}$, with $i=1,2$). The blue curve shows the numerical result for the single-mode case, where $\omega_{2}=g_{2}=\gamma_{2}=\gamma_{2m}=0$. (B and C) Experimental and numerical donor population evolution $P_{D}(t)$ versus energy gap $\Delta E$ and the number of vibrational oscillations of the fast mode $\omega_{1}t/2\pi$, using the same parameters as in the red circles and solid red curve in (A), respectively. Here, the detuning from the inner (outer) tilt mode, which encodes the fast (slow) vibrational energy, is set to $\delta_{1}=-2\pi\times 8$ kHz ($\delta_{2}=-2\pi\times 3$ kHz). Downward green arrows indicate nonadiabatic CT along $y_{1}$, assisted by single-phonon exchange with the slow mode via VAET.
Transitioning to VAET - To verify that the additional transfer pathways—provided by the slow vibrational mode—are responsible for the smooth transfer profile shown in Fig. 3D for the non-degenerate CT case, we decrease the electronic coupling strength and the vibronic coupling strength to the slow vibrational mode in Figs. 4A-C (see Appendix G for more details on VAET-CT crossover). Choosing a weaker electronic coupling ($|V|<\lambda_{1}/4$) while maintaining strong vibronic coupling to the fast mode ($g_{1}>\omega_{1}$) places the system closer to the nonadiabatic CT regime along the fast mode direction. In this case, the donor and acceptor populations remain localized on their respective potential energy surfaces along $y_{1}$, and resonant excitation transfers occur between well-defined donor and acceptor vibronic states at $\Delta E\approx\ell_{1}\omega_{1}$, with $\ell_{1}$ being an integer [26, 20] (the one-dimensional case of Eq. (2)). In the single-mode scenario, where only $y_{1}$ is considered, this results in the manifestation of the vibrational mode structure in the transfer rate spectrum (solid blue curve in Fig. 4A). Therefore, choosing a weaker electronic coupling allows us to distinguish the influences of the two vibrational degrees of freedom on the transfer dynamics when the second vibrational mode is involved. At the same time, lesser vibronic coupling to the slow mode reduces the distortion of the potential energy surfaces along $y_{2}$, making the transfer processes involving the slow mode to enter the VAET regime and introducing additional transfer rate resonances energetically enabled by the slow mode.
As shown by the resolved resonances in Fig. 4A, the additional peaks around the nonadiabatic transfer resonances of the fast mode, which coincide with the single-mode spectrum (solid blue curve), are induced by the second mode and located at $\Delta E\approx\sqrt{(\ell_{1}\omega_{1}+\ell_{2}\omega_{2})^{2}-(2V)^{2}}\approx\ell_{1}\omega_{1}+\ell_{2}\omega_{2}$, where the $\ell_{2}\omega_{2}$ contribution comes from the VAET process bridging the energy gap between vibronic states defined by the fast mode (see the discussion of the VAET regime below). The processes associated with $\ell_{2}=1$ dominate and are more clearly observable because of the perturbative nature of VAET. With either a larger electronic coupling strength $V$ or a stronger vibronic coupling strength to the slow mode $g_{2}$, the additional resonances caused by the presence of the slow mode broaden and merge with the fast-mode transfer resonances into a smooth spectrum, as shown, for example, in Fig. 3D. This highlights the crucial role of the simultaneous presence of fast and slow vibrational modes in reducing the CT dependence on the donor-acceptor energy gap through the additional transfer pathways.
Figure 5: Transfer rates in the VAET regime. (A) Transfer rate spectrum of degenerate VAET ($\omega_{1}\!=\!\omega_{2}\!\equiv\!\omega$) with $(V,g_{1},\gamma_{1},g_{2},\gamma_{2})\!=\!(0.200,0.200,0.045,0.220,0.018)\omega$. Red circles represent experimental data with error bars estimated via bootstrapping (see Methods). The solid red curve shows the transfer rate calculated from Eq. (4), using the definition in Eq. (5) and including spin decoherence ($\gamma_{z}=0.0007\omega$) and motional dephasing of both modes ($\gamma_{im}=0.0080\omega$, with $i=1,2$). The blue curve shows the numerical result of the single-mode VAET case, where $\omega_{2}\!=\!g_{2}\!=\!\gamma_{2}\!=\!\gamma_{2m}\!=\!0$. (B) Experimental and numerical donor population evolution $P_{D}(t)$ versus the number of vibrational oscillations $\omega t/2\pi$, using the same parameters as in the red circles and solid red curve in (A), respectively. Here, the detunings from the two tilt modes for encoding the degenerate vibrational energies are both set to $\delta_{1}=\delta_{2}=-2\pi\times 10$ kHz. (C and D) The same layout as in (A and B) for the non-degenerate VAET case ($\omega_{1}\!>\!\omega_{2}$) with $(V,g_{1},\gamma_{1},\omega_{2},\gamma_{z},\gamma_{im})=(0.167,0.275,0.028,0.667,0.0006,0.0068)\omega_{1}$ and $(g_{2},\gamma_{2})=(0.263,0.019)\omega_{2}$. The characteristic timescale is defined by the fast vibrational frequency $\omega_{1}$. In this case, the detuning from the inner (outer) tilt mode, which encodes the fast (slow) vibrational energy, is set to $\delta_{1}=-2\pi\times 12$ kHz ($\delta_{2}=-2\pi\times 8$ kHz). The red triangle, orange pentagon, green square, and blue circle markers correspond to $\Delta E=\{1.00,1.20,1.50,1.95\}\omega$ in the degenerate case and $\Delta E=\{0.75,0.92,1.42,1.60\}\omega_{1}$ in the non-degenerate case, respectively (marked by colored boxes in (A and C)). Near each resonant peak in the rate spectra, we label the total vibrational energy involved in the transfer.
@@
Vibrationally assisted exciton transfer (VAET) - Unlike in the CT regime, where strong vibronic coupling distorts the donor-acceptor potential energy landscape that defines the vibronic eigenstates of the system, in the VAET regime ($g_{i}\!\ll\!\omega_{i}$), the vibrational modes are weakly coupled to the electronic degree of freedom and therefore only act as facilitators of the exothermic transfer, where units of vibrational energy are exchanged with the electronic sites during the excitation transfer at specific $\Delta E$ resonances, given by Eq. (3) (see Fig. 5) [27]. Here, we also consider systems in the non-perturbative regime, characterized by strong electronic coupling ($|V|\sim\lambda_{i}/4$). For the vibrationally degenerate case ($\omega_{1}=\omega_{2}\equiv\omega$), shown in Fig. 5A, the transfer rate of the first resonance in the two-mode model is not enhanced compared to the single-mode case because only a single phonon from either degenerate mode can contribute to the transfer process at a time. In fact, the transfer rate becomes slightly slower due to the additional broadening from the dissipation of the second mode. On the contrary, at $\Delta E\approx\sqrt{(2\omega)^{2}-(2V)^{2}}$, VAET processes in which the total vibrational energy of $2\omega$ is supplied by a linear combination of the energy quanta from the two degenerate modes ($2\omega_{1}=2\omega_{2}=\omega_{1}+\omega_{2}$) interfere constructively. This leads to enhanced transfer rates compared to the single-mode case, where only the process involving a vibrational energy input of $2\omega_{1}$ can assist the transfer.
@@
In the non-degenerate case $(\omega_{1}>\omega_{2})$, the second-order processes occur at different donor-acceptor energy offsets, and their corresponding transfer resonances can be resolved experimentally (see Fig. 5C). As such, the presence of the slow mode does not affect the transfer rates at the resonances involving $\omega_{1}$ and $2\omega_{1}$ energy inputs. This is because, when we choose $\omega_{1}/\omega_{2}$ to be non-integer, there exists no non-trivial resonance involving an energy input of $\ell_{2}\omega_{2}$ from the slow mode, where $\ell_{1}\omega_{1}+\ell_{2}\omega_{2}=\ell\omega_{1}$. However, additional resonances emerge instead from processes involving energy inputs given by linear combinations of the two non-degenerate vibrational energies. In Fig. 5C, we observe three additional resonances beyond the two existing resonances provided by the fast vibrational mode for the two-mode case. These peaks correspond to the processes involving vibrational energy inputs of $\omega_{2}$, $2\omega_{2}$, and $\omega_{1}+\omega_{2}$.
@@
It is worth noting that, in both cases of mode degeneracy, the second-order processes associated with a vibrational energy contribution of $\omega_{1}+\omega_{2}$ consist of two pathways (one phonon from the fast mode, then another from the slow mode, and vice versa) that interfere constructively (see Fig. 1B). This coherent addition of the two rate amplitudes leads to an approximate two-fold enhancement of the peak at $\Delta E\approx 1.63\omega_{1}$ in Fig. 5C with respect to the neighboring peaks at $\Delta E\approx 1.29\omega_{1}$ and $\Delta E\approx 1.97\omega_{1}$ that are associated with the second-order processes involving $2\omega_{2}$ and $2\omega_{1}$ energy inputs, respectively (see Appendix E). This observation suggests that approximately half of the enhancement at $\Delta E\approx 1.95\omega$ in the two-mode degenerate case with respect to the single-mode case is attributed to the $\omega_{1}+\omega_{2}$ pathways (see Appendix E). While constructive interference of the two vibrational modes requires full coherent control over both degrees of freedom, our results also demonstrate that this coherent enhancement remains resilient in the presence of dissipation.
@@
In this work, we leverage the remarkable tunability of the trapped-ion platform to experimentally realize an open two-mode linear vibronic coupling model (LVCM) in two phenomenologically distinct regimes associated with charge transfer (CT) and vibrationally assisted exciton transfer (VAET), as well as their crossover. We simultaneously apply twelve carefully calibrated laser tones to independently control the coherent evolution of the qubit and the damping rates of two bosonic modes in a multi-species ion system. We observe enhanced transfer rates arising from the presence of the second mode across all vibronic coupling regimes—regardless of mode degeneracy—pointing out their differences and similarities. We also attribute these enhancements to coherent effects in the transfer pathways, which persist even in the presence of dissipation. Furthermore, our conclusions can be extended to two-site LVCM systems with more than two vibrational modes, where the same qualitative features identified in our findings remain (see Appendix F). This observation also highlights the necessity of considering anharmonicity in quantum dynamics models, where quantum scrambling can occur at resonances, transitioning coherent quantum behavior into chaotic dynamics [38].
@@
The experimental toolbox we deploy in this work is intrinsically scalable, as the same ion-laser couplings and sympathetic cooling techniques used to realize a two–mode, two–site LVCM can be extended to many vibrational modes and electronic sites without introducing additional physical overhead. In trapped-ion hardware, it is possible to include more molecular sites, as each extra qubit ion supplies a fully controllable two-level system that can encode a chromophore or charge-transfer center. Each additional qubit ion or coolant ion also adds three collective bosonic modes, which can be directly used for reservoir engineering by tailoring the spectrum of sympathetic-cooling lasers, without the need for digitization [39]. An arbitrary subset of these modes can then be endowed with individually programmable frequencies, coupling strengths, and dissipation rates, allowing the simulation of dissipative chemical dynamics in complex solvent environments [40, 41]. Employing multiple engineered bosonic modes also enables the experimental realization of spin-boson models with tunable spectral densities, formed by a linear superposition of Lorentzian components [42, 36, 21], crucial for exploring phenomena related to non-Markovian dynamics [43], such as coherence trapping [44, 45, 46, 47], and dissipative quantum state engineering [48].
@@ Methods
The experimental system used for this study is thoroughly described in Ref. [20], where the dynamics of the single-mode LVCM in the CT regime are realized with four Raman 355 nm and two 435.5 nm laser tones. In this work, we include two additional tones to each beam to generate the terms associated with the second vibrational mode, resulting in a total of ten laser tones on the 355 nm and 435.5 nm lasers, over which we have full control of both amplitudes and frequencies. We also use the collective modes of the two-ion chain along both radial directions of the trap—specifically the $y$ and $z$ tilt modes ($\omega_{\text{tilt},y}\equiv\omega_{\text{tilt},1}=2\pi\times 3.151$ MHz and $\omega_{\text{tilt},z}\equiv\omega_{\text{tilt},2}=2\pi\times 3.740$ MHz)—to encode the vibrational degrees of freedom in the two-mode LVCM Hamiltonian. The unused radial collective modes are the center-of-mass modes with $\omega_{\text{com},y}\equiv\omega_{\text{com},1}=2\pi\times 3.318$ MHz and $\omega_{\text{com},z}\equiv\omega_{\text{com},2}=2\pi\times 3.882$ MHz. Since the frequency separations among the available radial collective modes are set to be sufficiently large ($\Delta\omega_{\text{trap},j}\gtrsim 2\pi\times 140$ kHz), undesired off-resonant spin-phonon interactions, which are proportional to $g_{j}/(\mu_{i}-\omega_{\text{trap},j})$, can be neglected [23, 51]. Here, $\mu_{i}$ is the 355 nm laser frequency detuning from the qubit resonance, used to generate the spin-dependent force on the target radial tilt mode $i$, and $\omega_{\text{trap},j}$ is the relevant collective radial frequency considered for the off-resonant spin-phonon drive ($j\neq i$).
@@
[…]
@@
We experimentally simulate the LVCM dynamics from $t=0$ ms to $t=t_{\rm sim}$, where the finite simulation time $t_{\rm sim}$ ranges from 2 to 8 ms, corresponding to 24–65 vibrational cycles of the fast mode (defined by $\omega_{1}t/2\pi$), within which the system undergoing the first transfer resonance reaches equilibrium. Since Eq. (5) defines the transfer rate in the limit of $t_{\rm sim}\rightarrow\infty$, an offset correction is required when applying it to finite-time dynamics. Particularly, in the case of a strictly localized donor population, $P_{D}(t)={\rm constant}$, Eq. (5) still yields a nonzero transfer rate of $k_{T}=\frac{2}{t_{\rm sim}}$, which approaches zero only as $t_{\rm sim}\rightarrow\infty$. Therefore, we need to remove the undesired background from the finite-time evaluation of the non-zero steady-state donor population in the transfer rate calculations, as follows [20, 34]:
$k_{T}=\frac{\int_{0}^{t_{\text{sim}}}P_{D}(t)dt}{\int_{0}^{t_{\text{sim}}}tP_{D}(t)dt}-\frac{2}{t_{\text{sim}}}.$ (6)
We note that this background correction does not alter the characteristic features of the transfer rate spectra. By interpolating the donor population probability $P_{D}(t)$ for both the experimental and numerical data using the same time steps and applying the modified formula above, we obtain the transfer rates of the dynamics reported in the main text. As in Ref. [20], we also use a resampling (bootstrapping) procedure at each time step of the $P_{D}(t)$ measurements to estimate the uncertainties of the extracted transfer rates. For each time step, we treat the experimental error as the standard deviation of a normal distribution centered at the measured mean value, from which the resampled datasets are drawn. We then use the standard deviation of the transfer rates calculated from the resampled datasets as the uncertainty associated with the reported rates.